# SpiderCat: A Catalog of Compact Binary Millisecond Pulsars

### v1

@@ 3.5.2 Optical/IR
PSR J2339-0533 2440660623886405632 0.18 101333549113918682 0.17 233938.74-053305.2 0.20
Figure 8: The color-magnitude diagram showing extinction-corrected absolute g-band magnitudes versus g–r color for spider counterparts, overlaid on a background of three million Pan-STARRS field stars cross-matched with Gaia to obtain accurate distance and absolute magnitude measurements. Spider systems predominantly occupy a region between the main sequence and white dwarf branch. Outliers are, in most cases, either Huntsman systems (with giant companions), transitional MSPs (which may include optical emission from an accretion disk), or potentially unassociated stars (where the optical counterpart is uncertain due to a large angular separation from the pulsar or candidate). Most spider systems are located between the two dashed lines, where the upper boundary is defined by $M_{g}=4.5\times(g-r)+4.5$, and the lower by $M_{g}=2.7\times(g-r)+10.3$. We also draw two representative black body tracks with different temperatures from 3000 K to 16000 K separated by 1000 K (marked as stars) for BW-type (with radius $R_{\rm BW}=0.08\,R_{\odot}$) and RB-type stars (with radius $R_{\rm RB}=0.37\,R_{\odot}$).
@@
Unlike high-energy counterparts, the reliability of optical associations is more uncertain, particularly in regions of high stellar density. Therefore, we caution that not all associations may be genuine. Although we used a 2 arcsec search radius to accommodate high proper motions or positional uncertainties, separations exceeding $\sim$1 arcsec are questionable unless corroborated by other evidence (e.g., orbital period matching across wavebands). To highlight this, we mark in red the angular separations larger than 0.9 arcsec in Table 4.
We calculated the apparent and absolute optical magnitudes using Pan-STARRS and SkyMapper data, along with distances from Section 3.4. RB counterparts have extinction-corrected median g/r/i/z-band magnitudes of 18.5/18.3/18.1/17.8, with interquartile ranges of 17.3–19.4, 16.5–19.0, 16.2–18.6, and 16.0–18.5, respectively. BWs are on average $\sim$2 magnitudes fainter, with medians of 20.2/19.9/20.2/19.5 and interquartile ranges of 19.1–20.8, 19.0–21.1, 19.1–21.2, and 18.4–20.3. The median g–r color is 0.38 for RBs (interquartile range: 0.18–0.61) and 0.07 for BWs (interquartile range: -0.32–0.31). The individial values are reported in Table 5.
Figure 8 shows the color–magnitude diagram of spider systems, overlaid on a background of three million randomly selected Gaia stars with well-determined distances and magnitudes [^8] Selection based on DR3, following the criteria in Gaia Collaboration et al. (2018, Appendix B)., cross-matched with the Pan-STARRS catalog to provide consistent magnitudes and colors. Spider systems generally occupy a region between the main sequence and the white dwarf branch. Similar to what was done with Gaia data by J. Antoniadis (2021), we empirically define the region where the majority of spiders (75%) are found by drawing two lines: the upper boundary is given by $M_{g}=4.5\times(g-r)+4.5$, and the lower by $M_{g}=2.7\times(g-r)+10.3$. These lines serve as a rough guide for identifying potential companion candidates. However, for colors $(g-r)<0.25$, this linear approximation becomes less accurate, as the main sequence curves upward and spider systems appear above the defined region. Most outliers (10 out of 17 systems) are either huntsman systems with giant companions, transitional MSPs with possible optical emission from an accretion disk, or systems where the optical counterpart is uncertain due to a large angular separation from the pulsar or candidate coordinates. We note that the putative optical counterpart to the BW PSR J1928$+$1245 has an anomalously high $g$-band absolute magnitude of $M_{g}=0.7^{+1.1}_{-0.7}$ mag. This is the highest among confirmed spider systems and is approximately six magnitudes brighter than BWs at a similar color. Such discrepancy strongly suggests that this source is not the true counterpart, but rather an unrelated object along the line of sight to the pulsar, as already suggested by K. I. I. Koljonen & M. Linares (2023).
We also compute two representative color–magnitude tracks for BW- and RB-type companion stars, assuming blackbody spectra with temperatures ranging from 3000 K to 16000 K. These are shown in Fig. 8 as blue and orange lines, with star symbols marking blackbody temperatures in steps of 1000 K; cooler stars correspond to higher $(g-r)$ values. We adopt stellar radii by assuming the volume-equivalent Roche lobe radius (P. P. Eggleton, 1983), a neutron star mass of $1.8\,M_{\odot}$, mass ratios $q_{\rm BW}=0.01$ and $q_{\rm RB}=0.3$, median orbital periods of $P_{\rm orb,BW}=3.6$ hr and $P_{\rm orb,RB}=6.2$ hr (see Section 3.3), and a volume-equivalent Roche lobe filling factor of $f=0.6$. This results in $R_{\rm BW}=0.09\,R_{\odot}$ and $R_{\rm RB}=0.38\,R_{\odot}$ for BWs and RBs, respectively. Synthetic magnitudes are calculated using Pan-STARRS filter curves, assuming no extinction. These tracks agree quite well with the location of the majority of spider systems.
@@
Table 5: Spider distances and luminosities. See Sections 3.4, 3.5.1 and 3.5.2 for details.
Name d $F_{\rm X}$ $L_{\rm X}$ $F_{\rm\gamma}$ $L_{\rm\gamma}$ g $M_{g}$ g-r
(kpc) ($10^{-15}$ erg/s) ($10^{31}$ erg/s) ($10^{-12}$ erg/s) ($10^{33}$ erg/s) (mag) (mag) (mag)
PSR J0023+0923 $1.2^{+0.5}_{-0.5}$ $1.7\pm 0.3$ $0.3^{+0.4}_{-0.3}$ $7.8\pm 0.6$ $1.5^{+1.6}_{-1.0}$
PSR J0212+5321 $1.1^{+0.0}_{-0.0}$ $162.7\pm 18.4$ $24.5^{+4.2}_{-24.5}$ $15.6\pm 0.6$ $2.4^{+0.2}_{-0.2}$ 14.20 $3.9^{+0.1}_{-0.1}$ 0.24
PSR J0251+2606 $1.2^{+0.5}_{-0.5}$ $4.9\pm 0.4$ $0.8^{+0.9}_{-0.5}$
PSR J0312-0921 $0.8^{+0.3}_{-0.3}$ $5.5\pm 0.4$ $0.4^{+0.5}_{-0.3}$
4FGL J0327.3+2355 $1.8^{+0.9}_{-0.5}$ $2.2\pm 0.6$ $0.9^{+1.5}_{-0.5}$ 18.67 $7.4^{+0.7}_{-0.9}$ 0.43
4FGL J0336.0+7502 $2.4^{+1.9}_{-1.1}$ $2.1\pm 0.4$ $1.4^{+4.1}_{-1.4}$ $8.1\pm 0.5$ $5.4^{+13.3}_{-4.0}$ 20.46 $8.6^{+1.4}_{-1.3}$ 0.42
4FGL J0407.7-5702 $2.5^{+1.8}_{-0.9}$ $49.6\pm 1.3$ $37.0^{+76.8}_{-37.0}$ $1.6\pm 0.3$ $1.2^{+3.1}_{-0.8}$ 20.29 $8.3^{+1.0}_{-1.2}$ 0.38
3FGL J0427.9-6704 $2.5^{+0.5}_{-0.4}$ $117.6\pm 1.9$ $88.5^{+39.7}_{-88.5}$ $8.8\pm 0.5$ $6.6^{+3.3}_{-2.0}$ 17.89 $5.9^{+0.3}_{-0.4}$ 0.18
1FGL J0523.5-2529 $2.1^{+0.2}_{-0.2}$ $29.2\pm 3.0$ $15.4^{+4.9}_{-15.4}$ $12.0\pm 0.5$ $6.3^{+1.6}_{-1.2}$ 17.10 $5.5^{+0.2}_{-0.2}$ 0.65
PSR J0541+2959g
PSR J0610-2100 $3.3^{+1.3}_{-1.3}$ $7.2\pm 0.5$ $9.2^{+10.0}_{-6.1}$
PSR J0636+5128 $0.5^{+0.2}_{-0.2}$ $2.8\pm 0.4$ $0.1^{+0.1}_{-0.1}$
4FGL J0639.1-8009 $1.8^{+1.9}_{-0.9}$ $60.6\pm 6.1$ $22.9^{+85.8}_{-22.9}$ $2.5\pm 0.5$ $0.9^{+3.9}_{-0.7}$
4FGL J0705.8-0004 $2.4^{+1.8}_{-0.9}$ $3.2\pm 0.8$ $2.1^{+6.1}_{-1.5}$ 18.46 $6.6^{+1.0}_{-1.2}$ 0.62
3FGL J0737.2-3233 $0.7^{+0.0}_{-0.0}$ $10.2\pm 1.1$ $0.5^{+0.1}_{-0.1}$ 15.38 $6.3^{+0.1}_{-0.1}$ 0.34
PSR J0838-2827 $3.0^{+2.0}_{-1.2}$ $8.9\pm 0.6$ $9.5^{+18.6}_{-9.5}$ $7.3\pm 0.5$ $7.8^{+15.6}_{-5.2}$ 20.08 $7.7^{+1.1}_{-1.1}$ 0.62
2FGL J0846.0+2820 $3.7^{+0.6}_{-0.5}$ $1.1\pm 0.4$ $1.8^{+1.4}_{-1.8}$ 16.11 $3.2^{+0.3}_{-0.3}$ 0.57
4FGL J0935.3+0901 $1.2^{+1.1}_{-0.5}$ $13.8\pm 1.1$ $2.4^{+7.1}_{-2.4}$ $4.6\pm 0.5$ $0.8^{+2.5}_{-0.6}$ 21.07 $10.7^{+1.2}_{-1.4}$ 0.67
4FGL J0940.3-7610 $1.9^{+1.0}_{-0.5}$ $50.4\pm 15.7$ $21.0^{+45.8}_{-21.0}$ $6.2\pm 0.6$ $2.6^{+4.3}_{-1.4}$ 17.92 $6.6^{+0.7}_{-1.0}$ 0.61
PSR J0952-0607 $1.7^{+0.7}_{-0.7}$ $0.9\pm 0.2$ $0.3^{+0.5}_{-0.3}$ $2.4\pm 0.3$ $0.9^{+1.0}_{-0.6}$
PSR J0955-3947 $3.3^{+1.3}_{-1.3}$ $29.9\pm 6.1$ $39.6^{+53.8}_{-39.6}$ $11.6\pm 0.9$ $15.3^{+17.1}_{-10.2}$ 17.30 $4.7^{+1.1}_{-0.7}$0.21
PSR J1023+0038 $1.1^{+0.4}_{-0.4}$ $492.1\pm 1.2$ $72.9^{+70.3}_{-72.9}$ $32.4\pm 0.8$ $4.8^{+4.8}_{-3.1}$ 17.91 $7.7^{+1.1}_{-0.7}$ 1.37
PSR J1036-4353 $2.1^{+0.8}_{-0.8}$ $3.4\pm 0.5$ $1.8^{+2.3}_{-1.3}$ 19.63 $8.0^{+1.1}_{-0.7}$ 0.17
PSR J1048+2339 $2.0^{+0.8}_{-0.8}$ $8.7\pm 2.9$ $4.2^{+6.8}_{-4.2}$ $4.9\pm 0.5$ $2.3^{+2.7}_{-1.6}$ 19.97 $8.5^{+1.1}_{-0.7}$ 0.86
CXOU J110926.4-650224 $4.7^{+3.3}_{-2.2}$ $355.6\pm 2.7$ $947.8^{+1830.0}_{-947.8}$ $4.8\pm 1.0$ $12.8^{+32.5}_{-10.1}$
PSR J1124-3653 $1.0^{+0.4}_{-0.4}$ $6.0\pm 0.6$ $0.7^{+0.8}_{-0.7}$ $12.5\pm 0.6$ $1.5^{+1.5}_{-1.0}$
PSR J1221-0633 $1.3^{+0.5}_{-0.5}$ $5.8\pm 0.5$ $1.1^{+1.2}_{-0.7}$
PSR J1227-4853 $1.2^{+0.5}_{-0.5}$ $118.3\pm 1.1$ $21.9^{+21.4}_{-21.9}$ $19.0\pm 1.6$ $3.5^{+4.0}_{-2.4}$ 17.98 $7.5^{+1.1}_{-0.7}$ 0.48
PSR J1242-4712 $2.5^{+1.0}_{-1.0}$
PSR J1301+0833 $1.2^{+0.5}_{-0.5}$ $7.7\pm 0.5$ $1.4^{+1.5}_{-0.9}$
PSR J1302-3258 $1.4^{+0.6}_{-0.6}$ $0.9\pm 0.3$ $0.2^{+0.3}_{-0.2}$ $10.9\pm 0.5$ $2.7^{+2.8}_{-1.8}$
PSR J1306-4035 $1.4^{+0.6}_{-0.6}$ $55.6\pm 1.5$ $13.2^{+13.4}_{-13.2}$ $9.0\pm 0.7$ $2.1^{+2.4}_{-1.4}$ 18.39 $7.6^{+1.1}_{-0.7}$ 0.52
PSR J1311-3430 $2.4^{+1.0}_{-1.0}$ $17.6\pm 0.4$ $12.4^{+12.5}_{-12.4}$ $60.6\pm 1.2$ $42.6^{+42.6}_{-27.6}$ 18.62 $6.7^{+1.1}_{-0.7}$0.38
PSR J1317-0157 $2.8^{+1.1}_{-1.1}$ $1.5\pm 0.3$ $1.4^{+1.9}_{-1.0}$

### v4

@@ 3.5.2 Optical and Infrared
PSR J2339-0533 2440660623886405504 0.18 101333549113918682 0.17 233938.74-053305.2 0.20
Figure 8: Color-magnitude diagram showing extinction-corrected absolute $g$-band magnitudes vs. grg–r color for spider counterparts, overlaid on a background of three million Pan-STARRS field stars crossmatched with Gaia to obtain accurate distance and absolute magnitude measurements. Spider systems predominantly occupy a region between the main sequence and WD branch. Outliers are, in most cases, either HM systems (with giant companions), tMSPs (which may include optical emission from an accretion disk), or potentially unassociated stars (where the optical counterpart is uncertain due to a large angular separation from the pulsar or candidate). Most spider systems are located between the two dashed lines, where the upper boundary is defined by $M_{g}=4.5\times(g-r)+4.5$, and the lower by $M_{g}=2.7\times(g-r)+10.3$. We also draw two representative blackbody tracks with different temperatures from 3000 K to 16,000 K separated by 1000 K (marked as stars) for BW-type (with radius $R_{\rm BW}=0.09\,R_{\odot}$) and RB-type stars (with radius $R_{\rm RB}=0.38\,R_{\odot}$).
@@
The magnitudes reported from Pan-STARRS and SkyMapper are mean values from catalog photometry. Pan-STARRS typically provides 12 epochs per filter over a 5 yr survey, with single-epoch exposure times of 30–45 s, while SkyMapper typically provides six epochs per filter over a 7 yr survey, with single-epoch exposures of 100 s. Taken at random orbital phases, these observations generally average over the orbit and thus include flux contributions from both the intrinsic companion and the heated face. However, for faint sources, the detections may preferentially sample orbital phases near maximum brightness. We calculated the apparent and absolute optical magnitudes using Pan-STARRS and SkyMapper data, along with distances from Section 3.4. We corrected the optical magnitudes for extinction using reddening estimates from D. J. Schlegel et al. (1998) and extinction coefficients from E. F. Schlafly & D. P. Finkbeiner (2011) for Pan-STARRS, as well as those provided for SkyMapper on their website. [^9] https://skymapper.anu.edu.au/filter-transformations/ RB counterparts have extinction-corrected median $g$-, $r$-, $i$-, and $z$-band magnitudes of 18.5, 18.3, 18.1, and 17.8, with interquartile ranges of 17.3–19.4, 16.5–19.0, 16.2–18.6, and 16.0–18.5, respectively. BWs are on average $\sim$2 mag fainter, with medians of 20.2, 19.9, 20.2, and 19.5 and interquartile ranges of 19.1–20.8, 19.0–21.1, 19.1–21.2, and 18.4–20.3, respectively. The median grg–r color is 0.38 for RBs (interquartile range: 0.18–0.61) and 0.07 for BWs (interquartile range: -0.32–0.31). The individual values are reported in Table 5.
Figure 8 shows a color–magnitude diagram of spider systems, overlaid on a representative background of three million randomly selected Gaia stars with well-determined distances and magnitudes, [^10] Selection based on DR3, following the criteria in Gaia Collaboration et al. (2018, Appendix B). crossmatched with the Pan-STARRS catalog to provide consistent magnitudes and colors. Spider systems generally occupy a region between the main sequence and the WD branch. Similar to what was done with Gaia data by J. Antoniadis (2021), we empirically define the region where the majority of spiders (75%) are found by drawing two lines: the upper boundary is given by $M_{g}=4.5\times(g-r)+4.5$, and the lower by $M_{g}=2.7\times(g-r)+10.3$. These lines serve as a rough guide for identifying potential companion candidates. However, for colors (gr)<0.25(g-r)<0.25, this linear approximation becomes less accurate, as the main sequence curves upward and spider systems appear above the defined region. Most outliers (10 out of 17 systems) are either HM systems with giant companions, tMSPs with possible optical emission from an accretion disk, or systems where the optical counterpart is uncertain due to a large angular separation from the pulsar or candidate coordinates. We note that the putative optical counterpart to the BW PSR J1928++1245 has an anomalously high $g$-band absolute magnitude of $M_{g}=0.7^{+1.1}_{-0.7}$ mag. This is the highest among confirmed spider systems and is approximately 6 mag brighter than BWs at a similar color. Such discrepancy strongly suggests that this source is not the true counterpart, but rather an unrelated object along the line of sight to the pulsar, as already suggested by K. I. I. Koljonen & M. Linares (2023).
We also compute two representative color–magnitude tracks for BW- and RB-type companion stars, assuming blackbody spectra with temperatures ranging from 3000 K to 16,000 K. These are shown in Figure 8 as blue and orange lines, with star symbols marking blackbody temperatures in steps of 1000 K; cooler stars correspond to higher (gr)(g-r) values. We adopt stellar radii by assuming the volume-equivalent Roche-lobe radius (P. P. Eggleton, 1983), a neutron star mass of $1.8\,M_{\odot}$, mass ratios $q_{\rm BW}=0.01$ and $q_{\rm RB}=0.3$, median orbital periods of $P_{\rm b,BW}=3.6$ hr and $P_{\rm b,RB}=6.2$ hr (see Section 3.3), and a volume-equivalent Roche-lobe filling factor of $f=0.6$. This results in $R_{\rm BW}=0.09\,R_{\odot}$ and $R_{\rm RB}=0.38\,R_{\odot}$ for BWs and RBs, respectively. Synthetic magnitudes are calculated using Pan-STARRS filter curves, assuming no extinction. These tracks agree quite well with the location of the majority of spider systems.
[…]
Table 5: Spider Distances and Luminosities footnotetext: Notes. (This table is available in machine-readable form in the online article.)
Name d MethodaaWe estimate distances to spiders using dispersion measures (DMs) and the Galactic electron density model of J. M. Yao et al. (2017, “YMW16”), or, in some cases, the older model of J. M. Cordes & T. J. W. Lazio (2002, “NE2001”). When a DM is not available, we use geometric distances derived from Gaia Data Release 3 (DR3) parallaxes. However, if the relative parallax uncertainty is less than 0.2, we always adopt the parallax-based distance. The method used is noted in this column. See Section 3.4 for details. $F_{\rm X}$ Det.bbX-ray detectors and energy bands used to obtain the X-ray fluxes are as follows: XMM—EPIC on board XMM-Newton (0.2–12 keV); Swift—the X-ray Telescope (XRT) on board the Neil Gehrels Swift Observatory (0.3–10 keV); Chandra—ACIS on board the Chandra X-ray Observatory (0.2–7 keV); eRosita—eROSITA on board Spektrum-Roentgen-Gamma (0.2–8 keV). $L_{\rm X}$ $F_{\rm\gamma}$ $L_{\rm\gamma}$ $g$ $M_{g}$ grg-r
(kpc) ($10^{-14}$ ($10^{31}$ ($10^{-12}$ ($10^{33}$ (mag) (mag) (mag)
erg/s/cm2) erg/s) erg/s/cm2) erg/s)
PSR J0023+0923 $\sim$1.2 YMW16 $1.7\pm 0.3$ Chandra $0.3^{+0.4}_{-0.2}$ $7.8\pm 0.6$ $1.5^{+1.6}_{-1.0}$
PSR J0212+5321 $1.12^{+0.03}_{-0.03}$ Gaia $162.7\pm 18.4$ Swift $24.5^{+4.2}_{-3.8}$ $15.6\pm 0.6$ $2.4^{+0.2}_{-0.2}$ 14.20 $3.95^{+0.05}_{-0.06}$ 0.24
PSR J0251+2606 $\sim$1.2 YMW16 $4.9\pm 0.4$ $0.8^{+0.9}_{-0.5}$
PSR J0312-0921 $\sim$0.8 YMW16 $5.5\pm 0.4$ $0.4^{+0.5}_{-0.3}$
4FGL J0327.3+2355 $1.8^{+0.9}_{-0.5}$ Gaia $2.2\pm 0.6$ $0.9^{+1.5}_{-0.5}$ 18.67 $7.4^{+0.7}_{-0.9}$ 0.43
4FGL J0336.0+7502 $2.4^{+1.9}_{-1.1}$ Gaia $2.1\pm 0.4$ Chandra $1.4^{+4.1}_{-1.1}$ $8.1\pm 0.5$ $5.5^{+13.3}_{-4.0}$ 20.46 $8.6^{+1.4}_{-1.3}$ 0.42
4FGL J0407.7-5702 $2.5^{+1.8}_{-0.9}$ Gaia $49.6\pm 1.3$ XMM $37.4^{+75.7}_{-23.2}$ $1.6\pm 0.3$ $1.2^{+3.0}_{-0.8}$ 20.29 $8.3^{+1.0}_{-1.2}$ 0.38
3FGL J0427.9-6704 $2.5^{+0.5}_{-0.4}$ Gaia $117.6\pm 1.9$ XMM $8.4^{+40.1}_{-24.4} 8.8$$8\pm 0.5$ $6.6^{+3.3}_{-2.0}$ 17.89 $5.9^{+0.3}_{-0.4}$ 0.18
1FGL J0523.5-2529 $2.1^{+0.2}_{-0.2}$ Gaia $29.2\pm 3.0$ Swift $15.4^{+4.9}_{-3.7}$ $12.0\pm 0.5$ $6.3^{+1.6}_{-1.2}$ 17.10 $5.5^{+0.2}_{-0.2}$ 0.65
4FGL J0540.0-7552 $2.6^{+2.1}_{-1.1}$ Gaia $33.8\pm 1.0$ XMM $27.8^{+65.7}_{-18.1}$ $3.6\pm 0.5$ $3.0^{+8.1}_{-2.1}$
PSR J0541+2959g $\sim$1.5 YMW16
PSR J0610-2100 $\sim$3.3 YMW16 $7.2\pm 0.5$ $9.2^{+10.0}_{-6.1}$
PSR J0636+5128 $\sim$0.5 NE2001 $2.8\pm 0.4$ XMM $0.08^{+0.10}_{-0.06}$
4FGL J0639.1-8009 $1.8^{+1.9}_{-0.9}$ Gaia $60.6\pm 6.1$ XMM $23.0^{+86.7}_{-17.7}$ $2.5\pm 0.5$ $0.9^{+3.9}_{-0.7}$
4FGL J0705.8-0004 $2.4^{+1.8}_{-0.9}$ Gaia $3.2\pm 0.8$ $2.1^{+6.0}_{-1.5}$ 18.46 $6.6^{+1.0}_{-1.2}$ 0.62
3FGL J0737.2-3233 $0.66^{+0.02}_{-0.02}$ Gaia $10.2\pm 1.1$ $0.52^{+0.10}_{-0.08}$ 15.38 $6.29^{+0.07}_{-0.07}$ 0.34
PSR J0838-2827 $3.0^{+2.0}_{-1.2}$ Gaia $8.9\pm 0.6$ XMM $9.4^{+18.6}_{-6.2}$ $3\pm 0.5 7.7$$7^{+15.5}_{-5.1}$ 20.08 $7.7^{+1.1}_{-1.1}$ 0.62
2FGL J0846.0+2820 $3.7^{+0.6}_{-0.5}$ Gaia $1.1\pm 0.4$ Chandra $1.8^{+1.4}_{-0.9}$ 16.11 $3.2^{+0.3}_{-0.3}$ 0.57
4FGL J0935.3+0901 $1.2^{+1.1}_{-0.5}$ Gaia $13.8\pm 1.1$ XMM $2.4^{+7.1}_{-1.7}$ $4.6\pm 0.5$ $0.8^{+2.5}_{-0.6}$ 21.07 $10.7^{+1.2}_{-1.4}$ 0.67
4FGL J0940.3-7610 $1.9^{+1.0}_{-0.5}$ Gaia $50.4\pm 15.7$ Swift $21.0^{+46.2}_{-13.6}$ $6.2\pm 0.6$ $2.6^{+4.3}_{-1.4}$ 17.92 $6.6^{+0.7}_{-1.0}$ 0.61
PSR J0952-0607 $\sim$1.7 YMW16 $0.9\pm 0.2$ XMM $0.3^{+0.5}_{-0.2}$ $2.4\pm 0.3$ $0.9^{+1.0}_{-0.6}$
PSR J0955-3947 $\sim$3.3 NE2001 $29.9\pm 6.1$ Swift $39.6^{+53.8}_{-28.2}$ $11.6\pm 0.9$ $15.3^{+17.1}_{-10.2}$ 17.30 $4.7^{+1.1}_{-0.7}$ -0.21
PSR J1023+0038 $1.4^{+0.1}_{-0.1}$ Gaia $492.1\pm 1.2$ XMM $114.4^{+26.3}_{-19.6}$ $32.4\pm 0.8$ $7.5^{+1.9}_{-1.4}$ 17.91 $7.2^{+0.2}_{-0.2}$ 1.37
PSR J1036-4353 $\sim$2.1 NE2001 $3.4\pm 0.5$ $1.8^{+2.3}_{-1.3}$ 19.63 $8.0^{+1.1}_{-0.7}$ 0.17
PSR J1048+2339 $\sim$2.0 YMW16 $8.7\pm 2.9$ Swift $4.2^{+6.8}_{-3.2}$ $4.9\pm 0.5$ $2.3^{+2.7}_{-1.6}$ 19.97 $8.5^{+1.1}_{-0.7}$ 0.86
CXOU J110926.4-650224 $4.7^{+3.3}_{-2.2}$ Gaia $355.6\pm 2.7$ XMM $939.1^{+1806.5}_{-684.8}$ $4.8\pm 1.0$ $12.7^{+32.1}_{-10.0}$
PSR J1124-3653 $\sim$1.0 YMW16 $9.8\pm 3.3$ Swift $1.1^{+1.8}_{-0.9}$ $12.5\pm 0.6$ $1.5^{+1.5}_{-1.0}$
PSR J1221-0633 $\sim$1.3 YMW16 $5.8\pm 0.5$ $1.1^{+1.2}_{-0.7}$
PSR J1227-4853 $\sim$1.2 YMW16 $118.3\pm 1.1$ XMM $21.9^{+21.4}_{-14.1}$ $19.0\pm 1.6$ $3.5^{+4.0}_{-2.4}$ 17.98 $7.5^{+1.1}_{-0.7}$ 0.48
PSR J1242-4712 $\sim$2.5 YMW16
PSR J1301+0833 $\sim$1.2 YMW16 $5.7\pm 2.1$ Swift $1.0^{+1.7}_{-0.8}$ $7.7\pm 0.5$ $1.4^{+1.5}_{-0.9}$
PSR J1302-3258 $\sim$1.4 YMW16 $0.9\pm 0.3$ Chandra $0.2^{+0.3}_{-0.2}$ $10.9\pm 0.5$ $2.7^{+2.8}_{-1.8}$
PSR J1306-4035 $\sim$1.4 YMW16 $55.6\pm 1.5$ XMM $13.2^{+13.4}_{-8.6}$ $9.0\pm 0.7$ $2.1^{+2.4}_{-1.4}$ 18.39 $7.6^{+1.1}_{-0.7}$ 0.52
PSR J1311-3430 $\sim$2.4 YMW16 $17.6\pm 0.4$ XMM $12.4^{+12.5}_{-8.0}$ $60.6\pm 1.2$ $42.6^{+42.6}_{-27.6}$ 18.62 $6.7^{+1.1}_{-0.7}$ -0.38
PSR J1317-0157 $\sim$2.8 NE2001 $1.5\pm 0.3$ $1.4^{+1.9}_{-1.0}$
PSR J1356+0230 $\sim$1.8 YMW16 $1.9\pm 0.3$ $0.8^{+1.0}_{-0.5}$